# Bayesian Intrinsic Groupwise Image Registration: Unsupervised Disentanglement of Anatomy and Geometry

### v1

@@ 5.4 Evaluation Metrics
For experiments on the BraTS-2021 dataset, since the ground-truth spatial correspondence is available, the groupwise warping index (gWI) can be implemented as an additional metric, which measures the root mean squared error for unbiased deformation recovery based on the ground-truth and predicted deformation fields [6], i.e.
@@
The gWI provides a fine-grained measurement for voxel-wise co-registration accuracies, whereas the semantic evaluation metrics attend to high-level structural concurrence of distinct tissues. Besides, since all data have isotropic physical spacing, we measured the ASSD and gWI in voxel units for convenience.
@@ 5.5.1 Experimental Design
This experiment aims to showcase our model’s performance under various data conditions using the four datasets. These conditions include multi-modal data (all datasets except for OASIS), heterogeneous data such as BraTS-2021 of brain tumors and MS-CMRSeg of myocardium infarction, and intersubject groupwise registration (OASIS). Additionally, we aimed to demonstrate the efficiency of the partial learning strategy for our model. To this end, we examined four variants of our model, namely Ours-CN, Ours-CD, Ours-PN, and Ours-PD.
In this experiment, the test groups of each dataset maintained their original sizes, which was equal to the training group sizes used for both the baselines and complete learning of our model. However, in the case of partial learning for our model, the training group size was consistently set to 2, achieved by randomly selecting images before the experiment, rather than in each training iteration. This resulted in a reduced number of training images, posing additional challenges for model generalizability. Note that our model does not have a partial learning version for the Learn2Reg dataset (as each original group is an MR-CT pair) and the OASIS dataset (as it is mono-modal).

### v2

@@ 5.4 Evaluation Metrics
For experiments on the BraTS-2021 dataset, since the ground-truth spatial correspondence is available, the groupwise warping index (gWI) can be implemented as an additional metric, which measures the root mean squared error for unbiased deformation recovery based on the ground-truth and predicted deformation fields [6], i.e.,
@@
The gWI provides a fine-grained measurement for voxel-wise co-registration accuracies, whereas the semantic evaluation metrics attend to high-level structural concurrence of distinct tissues. Besides, since all data have isotropic physical spacing, we measured ASSD and gWI in voxel units for convenience.
@@ 5.5.1 Experimental Design
This experiment aims to showcase the performance of our model under various data conditions using the four datasets. These conditions include multi-modal data (all datasets except OASIS), heterogeneous data such as BraTS-2021 of brain tumours and MS-CMRSeg of myocardium infarction, and intersubject groupwise registration (OASIS). In addition, our objective was to demonstrate the efficiency of the partial learning strategy for our model. To this end, we examined four variants of our model, namely Ours-CN, Ours-CD, Ours-PN, and Ours-PD.
In this experiment, the test groups of each dataset maintained their original sizes, which was equal to the training group sizes used for both the baselines and complete learning of our model. However, in the case of partial learning for our model, the training group size was consistently set to 2, achieved by randomly selecting images before the experiment, rather than in each training iteration. This resulted in a reduced number of training images, posing additional challenges for model generalisability. Note that our model does not have a partial learning version for the Learn2Reg dataset (as each original group is an MR-CT pair) and the OASIS dataset (as it is mono-modal).